OpenStudy (anonymous):

Use the Pythagorean Theorem to determine if the three numbers could be the measures of the sides of a right triangle. Assume that the third measure is the hypotenuse. 12, 16, 20

OpenStudy (anonymous):

Can it be yes or no

OpenStudy (anonymous):

Plug the numbers into c^2 = a^2 + b^2 20^2 = 12^2 + 16^2 400 = 144 + 256 400 = 400 Since it is true, it is a right triangle.